Reports 1-1 of 1 Clear search Modify search
VAC (General)
shinji.miyoki - 7:59 Friday 25 November 2022 (23024) Print this report
Another leak in OMM tank

After today's Kimura-san's vacuum leak repair in the OMC and OMMT tank, TMP was successfully turned on (This report should be done by Kimura-san).

However, the vacuum level stopped at 0.12 Pa suddenly and kept constant. 

I re-entered the KAGRA center area and I confirmed that the TMP/Scroll pumps were working. So actually other leaks should exist.

I tried the "ethanol method", and found a leak at the GV-side d400 flange on the OMMT tank because the vacuum level enhanced from 0.12 Pa to 0.6 Pa when I put ethanol. Also I found several suspicious blank flanges and a connector flange on the OMMT and OMC tank.

I hope they will be repaired by Kimura-san.

Comments to this report:
nobuhiro.kimura - 7:48 Friday 25 November 2022 (23027) Print this report

When: 24/Nov./2022
Who: Kimura, Yasui and M. Takahashi

 We identified two leak points by spraying alcohol.  The leaks were in the K-400 flange joints with metal gaskets.

We retightened the flange joints at three locations, including these two, to stop the leaks.

For each K-400flange, 16 clamps were added and tightened to a torque of 60 Nm.

The remaining one flange was an ICF203 flange and was retightened with a torque of 15 Nm.

Search Help